Question: 70 students got an average score of 80, 30 students got an average score of 70. What is the average score of all students? Please explain how you calculate. thank you
Answer by stanbon(75887)

Question: 70 students got an average score of 80, 30 students got an average score of 70. What is the average score of all students? Please explain how you calculate.
-------------------
Average = total number of points/total number of students
----------------------
Total # of points = 70*80 + 30*70 =7700
Total # of students = 70 + 30 = 100
-----------------------------------------------
Average for all the students = 7700/100 = 77
===================================================
Cheers,
Stan H.